Form a polynomial whose zeros and degree are given.
Zeros: -7, multiplicity of 1; 2, multiplicity of 2; degree 3

Answer by ikleyn(52921)

.

p(x) = (x-(-7)*(x-2)^2 = (x+7)*(x-2)^2.